What are the prime factorizations of 33^4*39*4^14
I'm am completely confused on how to find this. Could someone please help me? It would be greatly appreciated.

Answer by jim_thompson5910(28717)

33 = 3*11

39 = 3*13

4 = 2*2

So

33^4*39*4^14

(3*11)^4*(3*13)*(2*2)^14

3^4*11^4*3*13*2^14*2^14

11^4*3^5*13*2^28

2^28*3^5*11^4*13


This means that

33^4*39*4^14 = 2^28*3^5*11^4*13